1/. A signal is transmitted from a ship and an echo is received from an object below the ship 4 s later. Calculate the distance of the object from the ship if the speed of sound in water is 1400 m/s.

1
Expert's answer
2022-04-11T12:44:48-0400
"2d=vt"

"d=\frac{vt}{2}=\frac{1400*4}{2}=2800\:\rm m"